OpenStudy (anonymous):

g(x) = x 2 – 1 Name: Transformation: helppp! i have no clue how to name a function or how to transform it!!! what do i do ???

OpenStudy (anonymous):

Did you mean to put g(x) = x^2 - 1? Right now the function doesn't really make sense.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Well the x^2 tells you it's a quadratic, right? x^1 is linear, x^2 is quadratic, x^3 is a cubic and so on. It's a quadratic function and notice the -1. That means the transformation is translated down 1 unit. In the same way, if that were x^2 + 1 it would make the function translated up 1 unit. Does that make sense?

OpenStudy (anonymous):

i understand the name part but not how you figure out the transformation part

OpenStudy (anonymous):

Ok. Let's call the transformation b. So the form of the function is g(x) = x^2 + b. If b = 0, we just have g(x) = x^2. Graphing that gives us a parabola that passes through (0, 0) Try it on your graphing calculator if you have one? Now if we have g(x) = x^2 + 3, that means the move the graph up 3 units but we don't move it to the right or left at all. That's the same way with your g(x) = x^2 - 1. We move the whole graph down 1 unit, but the shape and everything else stays the same.